# The Harrow-Hassidim-Lloyd algorithm with qutrits

### v1

@@ III.3 Qutrit HHL: working principles and quantum circuit
IV Application of qutrit HHL for molecular ground-state energy calculations
@@
In this section we discuss the application of qutrit HHL to quantum chemistry problems. For this purpose, we use the linearized form of coupled cluster equations [clarification] as our system of linear equations -
@@
We see that equation 1 takes the form of a system of linear equations, with the vector of cluster amplitudes being the unknown vector, $\vec{x}$.
The quantity of interest to us is the correlation energy which is given by -
@@
where $k=\parallel|x\rangle_{un}\parallel\;\parallel|{b}\rangle_{un}\parallel^{2}$. We note that we already know $|{b}\rangle_{un}$, while $|x\rangle_{un}$ can be obtained from $P(1)$ on the HHL ancilla qutrit. The molecular simulation problem is discussed in Reference [adaptHHL], one may refer to it and the references therein for detailed equations and derivations. Although we are interested in energy for the purposes of this work, one can in principle use the HHL algorithm for computing other molecular properties in the qubit or qutrit framework with an appropriate circuit module appended at the end of the HHL quantum circuit.
@@
V Results
V.1 Testing with toy matrices
@@
Table 1: Results from executing the qutrit HHL algorithm for the toy $A$ matrix and $\vec{b}$ described in equation (2), for different numbers of QPE clock register qubits. PFD refers to the percentage fraction difference between our result and the expected result obtained classically.
@@
respectively. Again, $\vec{b}$ is normalized for qutrit states, and can be directly encoded as $|b\rangle=|1\rangle$. This state can be achieved using a single qutrit increment gate. The statevector results obtained for this calculation using qutrit HHL for different values of $n_{r}$ are given in Table 2.
@@
V.2 Quantum chemistry simulations
We now move on to results from our simulations for quantum chemistry. We consider the $H_{2}$ molecule in the 631G6-31G basis with different bond lengths (9 data points between 1.2 and 1.6 Bohr) to obtain a potential energy curve (PEC). For each geometry considered, the system includes four molecular orbitals that comprise one occupied and three virtuals. The outermost virtual molecular orbital is frozen, hence our calculations consist of only three molecular orbitals for our active space that equals six spin orbitals. We use the GAMESS-US software [GAMESS] to generate the configuration interaction Hamiltonian in the singles and doubles approximation (CISD). In particular, we use Graphical unitary group approach (GUGA)-CISD so that the size of the Hamiltonian matrix reduces to $4\times 4$. Since the $A$ matrix obtained using linearized coupled cluster method in the singles and doubles approximation (LCCSD) is a principal sub-matrix of the Hamiltonian matrix, where we slice away the first row and first column (which is $\vec{b}$) of the latter [psiHHL], the state register in HHL has only one qutrit. We benchmark our results with energies obtained using the LCCSD approximation and CISD on a classical computer. On the HHL front, the one-qutrit isometry circuit for each bond length is a rotation gate, $R_{02}(\theta)$, with the choice of rotation angle depending on the bond length. Table A.1 of Appendix presents the values chosen for $\theta$.
The PEC is presented in Figure 5, and the accompanying data is presented in Table A.2 of the Appendix. Throughout, $n_{r}$ is set to 5 for our qutrit HHL calculations, while for the qubit HHL computations, we set $n_{r}=8$ to capture the results with comparable precision. However, for completeness, we also present the PEC with $n_{r}=5$ for qubit HHL in the figure as well as the table. We also add that many works that present quantum chemistry results in the quantum computing framework quote their methods’ energies relative to a benchmark value, but such a presentation does not provide complete context on how well an approach has captured correlation effects, given that one almost always starts with a reference wave function like the HF state to build the more accurate one and compute energies. Therefore, we also present the HF energies in the PEC as well as in the table of results, except that in Figure 5 we shift the HF energy at each bond length value, $R$, by a constant number for visual ease of discerning the total energy values provided using qubit HHL, qutrit HHL, LCCSD, and CISD methods. This is due to the fact that the correlation energy is typically much much smaller than the total energy or for that matter the HF energy, and therefore when we provide the HF energy in a figure along with the correlated values from methods like LCCSD, CISD, qubit HHL and qutrit HHL, one fails to discern the latter points. Our results show that as expected, qutrit HHL-LCCSD computations are able to capture correlation effects reasonably well, and the results are at most to within $0.02$ percent of the classical LCCSD value.

### v2

@@ III.3 Qutrit HHL: working principles and quantum circuit
IV Towards a practical implementation of the qutrit HHL
A very important part of a practical implementation of the qutrit HHL algorithm is the gadgets to use in decomposing the controlled-unitary gate that occurs in QPE. In fact, although our interest is in qutrit HHL, deriving such gadgets is useful for standalone Hamiltonian simulation and QPE calculations, as well as for other algorithms that use QPE. We feed in $A$ into the controlled-unitary, $C\;e^{iAt}$, and after expressing $A$ in the WH operator basis (Section IV.1) and Trotterizing (first order and first step), we arrive at a sequence of matrix exponentials each containing a WH operator in its exponent. Just as one uses Pauli gadgets to decompose each such exponential into a compact set of 1- and 2- qubit gates ($\sim n$ 2-qubit gates for an $n$-qubit circuit), we show in Section IV.2 that one can use WH gadgets for the qutrit counterpart. Different schemes are found in literature for decomposition of higher dimensional unitaries [shende_decompo, jiang_qutrit_decompo]. Recently, there has been a work on a decomposition similar to ours for multi-qutrit gates [decomposition_WHgadgets], where they extend the staircase structure of controlled-gates to the qutrit framework, with focus on decomposition using Gell-Mann matrices. We independently come up with Pauli gadget equivalents for qutrits, the WH gadgets, using the WH operators enlisted in Table 1. Figure 5 shows circuit diagrams for all possible WH gadgets for a 2-qutrit decomposition. Following is an outline of the stepwise implementation of qutrit HHL along with derivations wherever needed.
IV.1 Expressing the $A$ matrix in terms of WH operators
We can express any matrix of the form $2^{n}\times 2^{n}$ as a linear combination of tensor products of Pauli matrices. The coefficients of the expansion can be obtained using the relation [nielsenChuang] -
$A$ $=\sum_{k}c_{k}\,P_{k},$ (1)
$c_{k}$ $=\frac{1}{2^{n}}\operatorname{Tr}\!\left(P_{k}^{\dagger}A\right).$ (2)
Extending this relation to qutrits using the WH operators, for any matrix of dimension $3^{n}\times 3^{n}$ we get -
$A$ $=\sum_{a,b}c_{a,b}\,W_{a,b},$ (3)
$c_{a,b}$ $=\frac{1}{3^{n}}\operatorname{Tr}\!\left(W_{a,b}^{\dagger}A\right).$ (4)
Now, using Trotterization [trotter] (first order and Trotter step set to 1), we can approximate -
$e^{iAt}=e^{i\sum_{a,b}c_{a,b}\,W_{a,b}\,t}\approx\prod_{a,b}e^{i\,c_{a,b}\,W_{a,b}\,t}.$ (5)
IV.2 Circuit implementation of the unitary $e^{iAt}$ in terms of WH gadgets
Let us begin with deriving the 2-qutrit gadget for the simplest case of $Z\otimes Z$. Using the definition of $CX$ gate as -
$CX$ $=\sum_{j=0}^{2}|j\rangle\!\langle j|\otimes X^{j},$
and certain identities like $ZX=\omega XZ$, we can show that -
$CX^{\dagger}\,(I\otimes Z)\,CX$ $=\sum_{j=0}^{2}|j\rangle\!\langle j|\otimes X^{-j}ZX^{j}$
$=\sum_{j=0}^{2}\omega^{j}|j\rangle\!\langle j|\otimes Z$
$=Z\otimes Z.$
This can be extended to $Z^{2}$ which is equal to $Z^{\dagger}$ via -
$Z^{\eta}\otimes Z$ $=CX^{\dagger\eta}(I\otimes Z)CX^{\eta},\;\;\;\eta\in\{1,2\}.$ (6)
Now, using the well-known relation-
$e^{-i\Theta UPU^{\dagger}}$ $=Ue^{-i\Theta P}U^{\dagger},$ (7)
which is valid for any unitary $U$ and operator $P$, we can show that -
$e^{-\frac{i\Theta}{2}\left(Z^{\eta}\otimes Z+\text{h.c.}\right)}$ $=CX^{\dagger\eta}\;e^{-\frac{i\Theta}{2}\left(I\otimes Z+\text{h.c.}\right)}\;CX^{\eta}.$ (8)
The factor $\frac{1}{2}$ appears because of the Hermitian conjugate (h.c.) that is added to the WH string in order to preserve unitarity. The $e^{-\frac{i\Theta}{2}\left(I\otimes Z+\text{h.c.}\right)}$ can be implemented as a parametrized rotation about the z-axis for the second qutrit using $R_{Z}(\Theta)$ gate. The corresponding circuit is shown in Figure 5(a). In a similar way, the above logic can be extended to multiple qutrits via the staircase structure consisting of ${CX}^{\dagger\eta}$s to the left and ${CX}^{\eta}$s to the right, as shown in Figure 6.
We can further show that all the possible WH operators can be transformed to $Z^{2}$ or $Z$ via unitary operations. We derive these transformations for all the WH operators given in Table 1. Apart from the $Z$ and $Z^{2}$ operators, that can be directly implemented using Equation (8) we can show that $X=HZ^{2}H^{\dagger}$ and $X^{2}=HZH^{\dagger}$. The $XZ$ operator can be expressed as $SXS^{\dagger}$, where $X$ can be further broken down in terms of $H$ and $H^{\dagger}$, resulting in -
$XZ$ $=S(HZ^{2}H^{\dagger})S^{\dagger}.$ (9)
Now, on squaring Equation (9), we get -
$(XZ)^{2}$ $=(SHZ^{2}H^{\dagger}S^{\dagger})^{2}$
$XZXZ$ $=(SHZ^{2}H^{\dagger}S^{\dagger})(SHZ^{2}H^{\dagger}S^{\dagger})$
$\omega X^{2}Z^{2}$ $=SHZ^{4}H^{\dagger}S^{\dagger}$
$X^{2}Z^{2}$ $=SHZH^{\dagger}S^{\dagger},$
wherein we ignore the global phase factor $\omega$, or it can be equivalently absorbed in the coefficient in the exponent. We can also show that $Z=SZS^{\dagger}$, and using it along with Equation (9), we obtain -
$XZ^{2}$ $=(XZ)Z$
$=(SXS^{\dagger})(SZS^{\dagger})$
$=SXZS^{\dagger}$
$=S(S(HZ^{2}H^{\dagger})S^{\dagger})S^{\dagger},$
which can be further simplified as -
$XZ^{2}$ $=S^{\dagger}HZ^{2}H^{\dagger}S.$ (10)
The last WH operator that remains is $X^{2}Z$, which can be obtained by squaring Equation (10) as follows -
$(XZ^{2})^{2}$ $=XZ^{2}XZ^{2}$
$=\omega^{2}X^{2}Z^{4}$
$=\omega^{2}X^{2}Z.$
We can equate this to the right hand side of squared Equation (10) and ignore the global phase $\omega^{2}$. Therefore,
$X^{2}Z$ $=\left(S^{2}HZ^{2}H^{\dagger}S^{\dagger}\right)^{2}$ (11)
$=S^{\dagger}HZH^{\dagger}S.$
When any of the above WH operators occur on any one or both of the qutrits, they can be pulled outside the $CX^{\eta}$ and $CX^{\dagger\eta}$ occurring in Equation (8) using the relation in Equation (7). We show a simple demonstration of how the above-mentioned transformations translate into circuits for the case of $X$ occurring on the first qubit.
$e^{i\Theta(X\otimes Z)}$ $=e^{i\Theta(HZ^{2}H^{\dagger}\otimes Z)}$
$=e^{i\Theta\bigl((H\otimes I)(Z^{2}\otimes Z)(H^{\dagger}\otimes I)\bigr)}$
$=(H\otimes I)\,e^{i\Theta(Z^{2}\otimes Z)}\,(H^{\dagger}\otimes I).$
We can further expand $e^{i\Theta(Z^{2}\otimes Z)}$ using Equation (8). Circuit diagrams for the WH gadgets with all possible WH operators acting on the first qutrit are illustrated in Figure 5.
(a) The WH gadget for $e^{i\frac{\Theta}{2}(Z\otimes Z+h.c.)}$.
(b) The WH gadget for $e^{i\frac{\Theta}{2}(Z^{2}\otimes Z+h.c.)}$.
(c) The WH gadget for $e^{i\frac{\Theta}{2}(X\otimes Z+h.c.)}$.
(d) The WH gadget for $e^{i\frac{\Theta}{2}(X^{2}\otimes Z+h.c.)}$.
(e) The WH gadget for $e^{i\frac{\Theta}{2}(XZ\otimes Z+h.c.)}$.
(f) The WH gadget for $e^{i\frac{\Theta}{2}(X^{2}Z^{2}\otimes Z+h.c.)}$.
(g) The WH gadget for $e^{i\frac{\Theta}{2}(XZ^{2}\otimes Z+h.c.)}$.
(h) The WH gadget for $e^{i\frac{\Theta}{2}(X^{2}Z\otimes Z+h.c.)}$.
Figure 5: WH gadgets for all possible WH operators (excluding the trivial operator - Identity) acting on the first qutrit. Equivalent transformations can be applied to any other qutrit, for the respective WH operators acting on that qutrit.
Figure 6: The WH gadget for $e^{i\frac{\Theta}{2}(Z\otimes Z\otimes\ldots\otimes Z+h.c.)}$.
In this way, the expansion of a given AA matrix in terms of the WH operators (Equation (3)) can be implemented in a quantum circuit, as a unitary $U=e^{iAt}$ using Trotterization (Equation (5)).
IV.3 Efficient way of implementing the controlled-unitary in the qutrit framework
In the previous section, we saw how the unitary $U=e^{iAt}$ is implemented as a circuit. But our HHL implementation requires the controlled unitary $C\;e^{iAt}$. Instead of applying control over every gate in the gadget, we propose an efficient way of doing the same for qutrits. Reference [blunt_controlled_pauligadget] discusses an efficient way to decompose a controlled n-qubit gadget, into two (n+1)-qubit gadgets. In this section we extend this identity to the qutrit framework for the WH gadgets. The identity for Pauli gadgets in the qubit framework is given by -
$C\;U$ $=|0\rangle\!\langle 0|\otimes I+|1\rangle\!\langle 1|\otimes e^{i\Theta P_{j}}$
$=e^{\,i\Theta\left(|1\rangle\!\langle 1|\otimes P_{j}\right)}$
$=e^{\,\frac{i\Theta}{2}\left((I-Z)\otimes P_{j}\right)}$
$=e^{\,\frac{i\Theta}{2}\left(I\otimes P_{j}\right)}\;e^{-\frac{i\Theta}{2}\left(Z\otimes P_{j}\right)}.$
where $P_{j}$ is a Pauli operator acting on an n-qubit system. Consider the controlled-unitary for qutrits -
$U=|0\rangle\!\langle 0|\otimes I+|1\rangle\!\langle 1|\otimes e^{i\Theta W}+|2\rangle\!\langle 2|\otimes\left(e^{i\Theta W}\right)^{2},$ (12)
where $W=W_{a,b}$ are the WH operators used in Equation (3). We drop the subscript here for convenience. Let us define an operator -
$G=|1\rangle\!\langle 1|\otimes W+|2\rangle\!\langle 2|\otimes 2W.$ (13)
Using the orthogonality of the projectors $|0\rangle\!\langle 0|$, $|1\rangle\!\langle 1|$, and $|2\rangle\!\langle 2|$, and the fact that their powers give back the same projector, we have -
$G^{n}=|1\rangle\!\langle 1|\otimes W^{n}+|2\rangle\!\langle 2|\otimes(2W)^{n}.$ (14)
Consequently, the exponential of $G$ can be evaluated via its power series expansion -
$e^{i\Theta G}$ $=\sum_{n=0}^{\infty}\frac{(i\Theta)^{n}}{n!}G^{n}$
$=|0\rangle\!\langle 0|\otimes I+|1\rangle\!\langle 1|\otimes e^{i\Theta W}+|2\rangle\!\langle 2|\otimes e^{i\Theta 2W}.$ (15)
Now consider,
$C\;U=e^{i\Theta G}$ $=e^{\,i\Theta\left(|1\rangle\!\langle 1|\otimes W+|2\rangle\!\langle 2|\otimes 2W\right)},$ (16)
where we need to express the projectors as a linear combination of diagonal WH operators.
Expansion of diagonal operators in the Weyl–Heisenberg basis
For $\omega=e^{2\pi i/3}$, we can define the Z operator as -
$Z=\sum_{j=0}^{2}\omega^{j}|j\rangle\!\langle j|.$ (17)
The operators $\{I,Z,Z^{2}\}$ are diagonal in the computational basis and span the space of diagonal 1-qutrit operators.
Lemma 1.
Any diagonal operator $D$ acting on a single qutrit can be written as
$D=c_{0}I+c_{1}Z+c_{2}Z^{2},$ (18)
for suitable complex coefficients $c_{0},c_{1},c_{2}$.
Proof.
Since
$I$ $=\sum_{j=0}^{2}|j\rangle\!\langle j|,$ (19)
$Z$ $=\sum_{j=0}^{2}\omega^{j}|j\rangle\!\langle j|,$ (20)
$Z^{2}$ $=\sum_{j=0}^{2}\omega^{2j}|j\rangle\!\langle j|,$ (21)
the operators $\{I,Z,Z^{2}\}$ are linearly independent and form a basis for the three-dimensional space of diagonal operators. ∎
We now explicitly determine the expansion of the computational-basis projectors.
Lemma 2.
For $i\in\{0,1,2\}$, the rank-one projector $|i\rangle\!\langle i|$ admits the expansion
$|i\rangle\!\langle i|=\frac{1}{3}\left(I+\omega^{-i}Z+\omega^{-2i}Z^{2}\right).$ (22)
Proof.
Assume
$|i\rangle\!\langle i|=c_{0}I+c_{1}Z+c_{2}Z^{2}.$ (23)
Substituting the diagonal forms of $I$, $Z$, and $Z^{2}$ yields
$|i\rangle\!\langle i|=\sum_{j=0}^{2}\left(c_{0}+c_{1}\omega^{j}+c_{2}\omega^{2j}\right)|j\rangle\!\langle j|.$ (24)
Equating coefficients gives the linear system
$\delta_{ij}=c_{0}+c_{1}\omega^{j}+c_{2}\omega^{2j},\qquad j=0,1,2.$ (25)
In matrix form,
$\begin{pmatrix}\delta_{i0}\\ \delta_{i1}\\ \delta_{i2}\end{pmatrix}=\begin{pmatrix}1&1&1\\ 1&\omega&\omega^{2}\\ 1&\omega^{2}&\omega\end{pmatrix}\begin{pmatrix}c_{0}\\ c_{1}\\ c_{2}\end{pmatrix}$. (26)
The coefficient matrix is the discrete Fourier transform over $\mathbb{Z}_{3}$. Its inverse yields
$\begin{pmatrix}c_{0}\\ c_{1}\\ c_{2}\end{pmatrix}=\frac{1}{3}\begin{pmatrix}1&1&1\\ 1&\omega^{2}&\omega\\ 1&\omega&\omega^{2}\end{pmatrix}\begin{pmatrix}\delta_{i0}\\ \delta_{i1}\\ \delta_{i2}\end{pmatrix}$, (27)
from which the stated expression follows. ∎
From Lemma 2, each computational-basis projector admits the expansion
$|i\rangle\!\langle i|=\frac{1}{3}\left(I+\omega^{-i}Z+\omega^{-2i}Z^{2}\right),\qquad i\in\{0,1,2\}.$ (28)
The coefficients appearing in Equation (28) follow directly from the inverse discrete Fourier transform over $\mathbb{Z}_{3}$ and are uniquely determined by the diagonal eigenvalues of the WH operator, $Z$.
Substituting $i=0$ into Equation (28) yields
$|0\rangle\!\langle 0|$ $=\frac{1}{3}\left(I+Z+Z^{2}\right).$ (29)
For $i=1$, using $\omega^{-1}=\omega^{2}$ and $\omega^{-2}=\omega$, we obtain
$|1\rangle\!\langle 1|$ $=\frac{1}{3}\left(I+\omega^{2}Z+\omega Z^{2}\right).$ (30)
Finally, for $i=2$, noting that $\omega^{-2}=\omega$ and $\omega^{-4}=\omega^{2}$, we find
$|2\rangle\!\langle 2|$ $=\frac{1}{3}\left(I+\omega Z+\omega^{2}Z^{2}\right).$ (31)
Substituting Equations (30) and (31) in Equation (16) we get -
$C\;U$ $=e^{\,i\Theta\left(|1\rangle\!\langle 1|\otimes W+|2\rangle\!\langle 2|\otimes 2W\right)}$
$=e^{\,i\Theta\left(\frac{1}{3}\left(I+\omega^{2}Z+\omega Z^{2}\right)\otimes W+\frac{1}{3}\left(I+\omega Z+\omega^{2}Z^{2}\right)\otimes 2W\right)}$
$=e^{\,\frac{i\Theta}{3}\Bigl((3I+(2\omega+\omega^{2})Z+(\omega+2\omega^{2})Z^{2})\otimes W\Bigr)}$
which can be further simplified as -
$C\;U$ $=\left(e^{\,i\Theta(I\otimes W)}\right)\left(e^{\,\frac{i\Theta(2\omega+\omega^{2})}{3}(Z\otimes W)}\right)\left(e^{\,\frac{i\Theta(\omega+2\omega^{2})}{3}(Z^{2}\otimes W)}\right).$ (32)
Thus, we obtain a product of three WH gadgets on simplification of the controlled-unitary, which is a more efficient way to implement it in a quantum circuit.
V Application of qutrit HHL for molecular ground-state energy calculations
@@
In this section we discuss the application of qutrit HHL to quantum chemistry problems. For this purpose, we use the linearized form of coupled cluster equations as our system of linear equations -
@@
We see that Equation (33) takes the form of a system of linear equations, with the vector of cluster amplitudes being the unknown vector, $\vec{x}$.
The quantity of interest to us is the correlation energy, which is given by -
@@
where $k=\parallel|x\rangle_{un}\parallel\;\parallel|{b}\rangle_{un}\parallel^{2}$. We note that we already know $|{b}\rangle_{un}$, while $\parallel|x\rangle_{un}\parallel$ can be obtained from $P(1)$ on the HHL ancilla qutrit. The molecular simulation problem is discussed in Reference [adaptHHL], one may refer to it and the references therein for detailed equations and derivations. Although we are interested in energy for the purposes of this work, one can in principle use the HHL algorithm for computing other molecular properties in the qubit or qutrit framework with an appropriate circuit module appended at the end of the HHL quantum circuit.
@@
VI Results
VI.1 Testing with toy matrices
@@
Table 2: Results from executing the qutrit HHL algorithm for the toy $A$ matrix and $\vec{b}$ described in Equation (34), for different numbers of QPE clock register qutrits. PFD refers to the percentage fraction difference between our result and the expected result obtained classically.
@@
respectively. Again, $\vec{b}$ is normalized for qutrit states, and can be directly encoded as $|b\rangle=|1\rangle$. This state can be achieved using a 1-qutrit increment gate. The statevector results obtained for this calculation using qutrit HHL for different values of $n_{r}$ are given in Table 3.
@@
VI.2 Quantum chemistry simulations
We now move to results from our simulations for quantum chemistry. We use the GAMESS-US software [GAMESS] to generate the Graphical Unitary Group Approach Configuration Interaction Hamiltonian in the Singles and Doubles approximation (GUGA-CISD). Since the $A$ matrix is obtained using linearized coupled cluster method in the singles and doubles approximation (LCCSD), it is a principal sub-matrix of the Hamiltonian matrix, where we slice away the first row and first column (which is $\vec{b}$) of the latter [psiHHL]. We benchmark our results with energies obtained using the LCCSD approximation and CISD on a classical computer.
Case 1: 1-qutrit input state
We consider the $H_{2}$ molecule in the 631G6-31G basis with different bond lengths (9 data points between 1.2 and 1.6 Bohr) to obtain a potential energy curve (PEC). For each geometry considered, the system includes four molecular orbitals that comprise one occupied and three virtuals. The outermost virtual molecular orbital is cut-off, hence our calculations consist of only three molecular orbitals for our active space that equals six spin orbitals, thus yielding a $3\times 3$ AA matrix in the GUGA-CISD framework. On the HHL front, the 1-qutrit isometry circuit for each bond length is a rotation gate, $R_{02}(\theta)$, with the choice of rotation angle depending on the bond length. Table A.1 of Appendix presents the values chosen for $\theta$.
The PEC is presented in Figure 7, and the accompanying data is presented in Table A.2 of the Appendix. Throughout, $n_{r}$ is set to 5 for our qutrit HHL calculations, while for the qubit HHL computations, we set $n_{r}=8$ to capture the results with comparable precision. However, for completeness, we also present the PEC with $n_{r}=5$ for qubit HHL in the figure as well as the table. We also add that many works that present quantum chemistry results in the quantum computing framework quote their methods’ energies relative to a benchmark value, but such a presentation does not provide complete context on how well an approach has captured correlation effects, given that one almost always starts with a reference wave function like the HF state to build the more accurate one and compute energies. Therefore, we also present the HF energies in the PEC as well as in the table of results, except that in Figure 7 we shift the HF energy at each bond length value, $R$, by a constant number for visual ease of discerning the total energy values provided using qubit HHL, qutrit HHL, LCCSD, and CISD methods. Our results show that as expected, qutrit HHL-LCCSD computations are able to capture correlation effects reasonably well, and the results are at most to within $0.02$ percent of the classical LCCSD value.